Consider a 0.52 g puck that moves as shown in the
Answer #1

From graph, its clear that relation between x and y coordinate is

y=5x

Since 60 pictures were taken per second hence time interval between two consecutive picture will be 1/60 sec. i.e.

\Delta t=\frac{1}{60}

In this same time interval, x coordinate change by 0.1 cm while y coordinate change by 0.5 cm i.e.

\Delta x=0.001 m\\ \Delta y=0.005 m

(a) Velocities

v_x=\frac{\Delta x}{\Delta t}=\frac{.001}{1/60}=0.06 m/sec\\ v_y=\frac{\Delta y}{\Delta t}=\frac{.005}{1/60}=0.30 m/sec

(b) Velocities in terms of i and j

V=v_x\hat{i}+v_y\hat{j}=(0.06\hat{i}+0.3\hat{j})m/sec

Magnitude of velocity

V=|V|=|v_x\hat{i}+v_y\hat{j}|=(\sqrt{0.06^2+0.3^2})m/sec=0.3059 m/sec

In terms of magnitude and angle

\tan{\theta}=\frac{v_y}{v_x}\Rightarrow \theta=tan^{-1}(\frac{v_y}{v_x})=tan^{-1}(\frac{0.3}{0.06})=78.7^o

therefore

\vec{V}=V\cos{\theta}\hat{i}+V\sin{\theta}\hat{j}=0.3059\cos{(78.7)}\hat{i}+0.3059\sin{(78.7)}\hat{j}

(c) Momentum in i and j

\vec{P}=m(v_x\hat{i}+v_y\hat{j})=0.52\times10^{-3}\times(0.06\hat{i}+0.3\hat{j})kgm/sec\\ =(3.12\hat{i}+15.6\hat{j})kgm/sec

In terms of magnitude and angle

\vec{P}=m(V\cos{\theta}\hat{i}+V\sin{\theta}\hat{j})=0.52\times 10^{-3}\times(0.3059\cos{(78.7)}\hat{i}+0.3059\sin{(78.7)}\hat{j})\\ =0.0063\cos{(78.7)}\hat{i}+0.0063sin{(78.7)}\hat{j}